OpenStudy (anonymous):

Which phrase represents the algebraic expression 3 + x/5 ? A. five divided by a number, increased by three B. five greater than a number divided into three C. the sum of three and a number divided by five D. five more than the quotient of a number and three

jimthompson5910 (jim_thompson5910):

is it \[\large \frac{3+x}{5}\] or \[\large 3+\frac{x}{5}\] ??

OpenStudy (anonymous):

I'd go with C

OpenStudy (anonymous):

@jim_thompson5910 the second one !

jimthompson5910 (jim_thompson5910):

so yes, it would be C since 3 + (x/5) = sum of 3 and x/5 3 + (x/5) = sum of three and a number divided by 5

OpenStudy (anonymous):

thank you :) ! <3

jimthompson5910 (jim_thompson5910):

yw